OpenStudy (anonymous):

how to integrate the following: Integration [(1+cos x ) /sin x . cos x ] w. r. t. dx.

OpenStudy (anonymous):

anyone please??/

OpenStudy (anonymous):

are w. r. t constants? if so. 1+cos x/(sin x . cos x) = 1/(sin x . cos x) + 1/(sin x) = 2/sin 2x + 1/(sin x) = 2 . Cosec 2x + Cosec x. in this form you can find the integral in an integration table. I hope it helps.

OpenStudy (anonymous):

yep.. thank you sir @juancarlosquintero_EC

OpenStudy (anonymous):

Take 1+cosx = cos^2(x/2)+sin^2(x/2)+cos^2(x/2)-sin^2(x/2) = 2*cos^2(x/2) Similarly, break the terms in denominator and then cancel the common terms. Finally write dx=2*d(x/2) And integrate the expression wrt d(x/2). :)